# Clustering-Based Sub-QUBO Extraction for Hybrid QUBO Solvers

### v1

@@ Abstract
Quantum Approximate Optimization Algorithm (QAOA) can be used to solve quadratic unconstrained binary optimization (QUBO) problems. However, the size of the solvable problem is limited by the number of qubits. To leverage noisy intermediate-scale quantum (NISQ) devices to solve large-scale QUBO problems, one possible way is to decompose the full problem into multiple sub-problems, which we refer to as the Sub-QUBO Formalism. In this work, we enhance this formalism by proposing a sub-QUBO extraction protocol. To do so, we define a measure to quantify correlations between variables and use it to build a correlation matrix. This matrix serves as the input for clustering algorithms to group variables. Variables belonging to the same group form sub-QUBOs and are subsequently solved using QAOA. Our numerical analysis on several classes of randomly generated QUBO problems demonstrates that this grouping method outperforms previous approaches in terms of objective function values, while maintaining a comparable number of quantum subroutine calls. This method offers wide applicability for solving QUBO problems on NISQ devices.
@@ I Introduction
Quadratic unconstrained binary optimization (QUBO) is a fundamental NP-hard problem in combinatorial optimization, with significant applications in fields such as finance, economics, logistics, production planning, and machine learning [1, 2]. Although it is widely accepted that exactly solving general QUBO problems requires exponential time [3], various classical heuristic algorithms can approximate the solutions of QUBO problems. These algorithms have been extensively studied, particularly those based on local search methods, such as simulated annealing [4, 5] and tabu search [6, 7]. Population-based algorithms are also often used to solve QUBO problems, such as genetic algorithms [8, 9].
The Quantum Approximate Optimization Algorithm (QAOA) provides a quantum approach to solve these complex optimization problems [10, 11, 12]. QAOA utilizes quantum computers by encoding the QUBO problem into a cost Hamiltonian and applying alternating layers of time evolution governed by this Hamiltonian and a transverse field. By optimizing $O(p)$ parameters, where $p$ denotes the number of layers, the output state converges to an approximate solution [10, 13].
However, the number and connectivity of qubits in NISQ devices pose significant challenges to the application of QAOA on solving QUBO problems [12, 14]. To overcome these limitations, researchers have applied the concept of problem decomposition to QAOA, resulting in the development of hybrid QUBO solvers with sub-QUBO extraction [15, 16]. This hybrid approach, which combines QAOA with classical search, enables exploration of a broader solution space and reduces the likelihood of the algorithm being trapped in local optima while addressing the constraints of qubit counts and connectivity. For clarity, we refer to such algorithms as the sub-QUBO formalism in the following text.
@@
In the sub-QUBO formalism, the main QUBO objective is iteratively decomposed into smaller sub-QUBOs by grouping variables based on the QUBO matrix and the current solution. Each sub-QUBO problem is solved using QAOA, and their solutions are combined and refined to a local minimum via classical optimization. The resulting refined solution is used as the initial guess for the next iteration. This iterative process continues until convergence of the objective function is reached. A critical component of the sub-QUBO formalism is the grouping strategy, as the quality of grouping directly impacts both the quality of the final solution and the convergence rate of the algorithm. Various strategies have been proposed for variable partitioning, including random selection [15], sorting variables based on the impact of each variable on the objective function (termed Impact-Indexing)[16], and extracting the most undetermined variables from a pool of previous quasi-solutions (termed Certainty-Degree)[15, 17].
@@
In this study, we propose a clustering-based variable grouping scheme in sub-QUBO formalism. Firstly, we construct a correlation matrix to capture the joint contributions of variables to the objective function. The correlation matrix can be derived from the QUBO matrix and the current solution. Then, we employ the multi-view spectral clustering method [18, 19] to group variables into clusters according to their correlations. This clustering reduces inter-cluster interactions, leading to enhanced solution quality. We evaluate the performance of our algorithm on Max-Cut problems of randomly generated regular graphs and Erdős–Rényi graphs. Through numerical simulations, we demonstrate that our method consistently achieves lower objective function values compared to existing approaches. Additionally, our approach requires a comparable number of quantum subroutine calls to previous methods and even fewer in cases with larger sub-QUBO sizes.
@@ II.1 QUBO Problems and QAOA
where $\mathbf{x}$ is a vector of binary variables, $Q$ is a square matrix representing the coefficients of the quadratic terms, and $\mathbf{c}$ is a vector containing the coefficients of the linear terms. Given the binary nature of the variables ($x_{i}\in\{0,1\}$), the linear terms can be equivalently transformed into quadratic terms, enabling the entire problem to be reformulated in a purely quadratic form. The significance of the QUBO framework lies in its ability to encompass a broad range of optimization problems. Most of COPs can be reformulated by encoding constraints and objectives into binary variables and a quadratic cost matrix [20].
@@
Solving the ground state energy of the Ising model is a NP-hard problem. Classical heuristic algorithms can provide approximate solutions [24]. Specialized computers, such as coherent Ising machines, offer alternative methods that can yield accurate approximations at high speeds [25]. Additionally, quantum computing provides another promising approach for tackling Ising problems, as demonstrated in recent experiments using both analog [26, 27, 28, 29, 30] and digital quantum simulations [31, 32, 33, 34].
After mapping the QUBO problem to the Ising Hamiltonian, QAOA can be employed to approximate the ground state of the Hamiltonian, which corresponds to the minimum of the objective function. QAOA [10, 11] is a variational quantum algorithm that approximates the ground state of a Hamiltonian using parameterized quantum circuits. As illustrated in Fig. 1(f), the QAOA involves two key components: the cost Hamiltonian $H_{C}$, which corresponds to the Ising Hamiltonian in our case, and the mixer Hamiltonian $H_{M}=\sum_{i}\sigma_{x}^{i}$ that facilitates exploration of the solution space.
@@ II.2 The Sub-QUBO Formalism
In this formalism, three components allow flexible implementation choices. The first is the grouping method, discussed in detail in Sec. II.3. The second component is the sub-QUBO solver, which can be quantum annealing, QAOA, or even classical algorithms. In this work, we focus primarily on QAOA. Finally, the classical search that forces the solution to a local minimum can also switch between multiple choices. In our work, we employ a simple greedy local search to emphasize the role of the quantum subroutine, while previous studies have explored other classical search methods, including the tabu search [7, 35] as used in Ref. [16]. A discussion on the performance of tabu search within this formalism is provided in appendix A.
@@ II.3 Grouping Methods in the Sub-QUBO Formalism
The effectiveness of hybrid annealing algorithms critically depends on how sub-QUBOs are extracted. Ref. [15] introduced the hybrid annealing framework and several grouping strategies, including random selection and an Impact-Indexing approach (originally referred to as gains-based). The Impact Indexing method measures the ’impact’ of each variable by evaluating how much the objective function changes when that variable is flipped. Initially, Ref. [15] formed sub-QUBOs by choosing the variables with the largest negative impacts; however, this often led to suboptimal solutions trapped in local minima.
@@ IV.1 Dataset
IV.2 Performance Evaluation
@@
Figure 4: Performance comparison and computational resource requirements for three sub-QUBO algorithms solving Max-Cut problems with sub-problem sizes ranging from 10 to 24 qubits. The algorithms use Impact-Indexing (blue), Certainty-Degree (green), and clustering (red) as grouping methods. (a,d) Minimum objective function values for 100 randomly generated Max-Cut problems based on (a) 100-node 3-regular graphs and (d) 100-node Erdős–Rényi graphs. Lower values indicate better performance. (b,e) Number of quantum QAOA subroutine calls versus sub-problem size for (b) 3-regular graphs and (e) Erdős–Rényi graphs. (c,f) Total number of QAOA iterations versus sub-problem size for (c) 3-regular graphs and (f) Erdős–Rényi graphs. Shaded areas represent 95% confidence intervals.
@@
In Fig. 4(b) and (e), we observe a sharp decrease in the number of quantum subroutine calls when the sub-QUBO size increases from $15$ to $16$ for the Certainty-Degree and Impact-Indexing grouping methods. Further numerical experiments on varying Max-Cut problem sizes using three grouping methods reveal that both Certainty-Degree and Impact-Indexing methods exhibit sudden drops in the number of subroutine calls at specific ratios of sub-QUBO size to problem size. In contrast, the clustering method demonstrates a more consistent behavior, with the number of calls approximately inversely proportional to the sub-QUBO size across all problem sizes. Detailed discussions of these results are conducted in appendix B.
@@ Appendix B Algorithm Complexity Analysis
Without dividing into sub-problems and under certain assumptions, such as parallel implementation of time evolution under $H_{C}$, the time complexity of QAOA for solving the Max-Cut problem is primarily determined by the ansatz depth $p$ rather than the problem size $N$ [53, 54]. Specifically, for a $D$-regular graph with finite $D$, the time complexity of QAOA is $O(p16^{p})$. For a fixed $p$, the time complexity of the quantum subroutine can be considered constant. In the context of the sub-QUBO formalism, the overall time complexity can be effectively represented by the number of quantum subroutine calls. This section analyzes how the number of quantum subroutine calls varies with the problem size $N$ and the sub-QUBO size $d$ under different grouping methods.

### v2

@@ Abstract
Quantum Approximate Optimization Algorithm (QAOA) offers a promising approach for solving quadratic unconstrained binary optimization (QUBO) problems. However, the size of the solvable problem is limited by the number of qubits. To leverage noisy intermediate-scale quantum (NISQ) devices to solve large-scale QUBO problems, one possible way is to decompose the full problem into multiple sub-problems, which we refer to as the Sub-QUBO Formalism. In this work, we enhance this formalism by proposing a sub-QUBO extraction protocol. To do so, we define a measure to quantify correlations between variables and use it to build a correlation matrix. This matrix serves as the input for clustering algorithms to group variables. Variables belonging to the same group form sub-QUBOs and are subsequently solved using QAOA. Our numerical analysis on several classes of randomly generated QUBO problems demonstrates that this grouping method outperforms previous approaches in terms of objective function values, while maintaining a comparable number of quantum subroutine calls. This method offers wide applicability for solving QUBO problems on NISQ devices.
@@ I Introduction
Quadratic unconstrained binary optimization (QUBO) is a fundamental NP-hard problem in combinatorial optimization [1, 2]. Various classical heuristic algorithms can approximate the solutions of QUBO problems. These algorithms have been extensively studied, particularly those based on local search methods, such as simulated annealing [3, 4] and tabu search [5, 6].
In the quantum domain, solving QUBO is equivalent to finding the ground state of an Ising Hamiltonian. Several quantum algorithms have been developed for this purpose. Quantum Signal Processing (QSP) filter algorithms [7] and Lindbladian-based approaches [8, 9] require fault-tolerant quantum computers to find low-energy states. Algorithms based on the Hadamard test and classical Fourier transform are applicable to early fault-tolerant quantum computers [10], while quantum annealing requires specialized hardware [11]. Variational quantum algorithms, such as the Quantum Approximate Optimization Algorithm (QAOA), can be implemented on near-term intermediate-scale quantum (NISQ) devices [12, 13, 14]. QAOA encodes the QUBO problem into a cost Hamiltonian and applies alternating layers of time evolution governed by this Hamiltonian and a transverse field. By optimizing $O(p)$ parameters, where $p$ denotes the number of layers, the algorithm aims to approximate the ground state solution [12, 15].
However, the practical deployment of QAOA on current NISQ devices is constrained by the limited number of high-fidelity qubits, shallow circuit depths, and noise accumulation, which restrict its applicability to small-scale problems [14, 16, 17]. To overcome these limitations, researchers have applied the concept of problem decomposition to QAOA, resulting in the development of hybrid QUBO solvers with sub-QUBO extraction [18, 19, 20, 21, 22, 23].
@@
Within the sub-QUBO formalism, the main QUBO objective is iteratively decomposed into smaller sub-QUBOs by grouping variables. Each sub-QUBO problem is solved using a quantum subroutine, and their solutions are combined and refined to a local minimum via classical optimization. A critical component of the sub-QUBO formalism is the grouping strategy, as the quality of grouping directly impacts both the quality of the final solution and the convergence rate of the algorithm. Various strategies have been proposed for variable partitioning, including random selection [18], sorting variables based on the impact of each variable on the objective function (termed Impact-Indexing)[19], and extracting the most undetermined variables from a pool of previous quasi-solutions (termed Certainty-Degree)[18, 24].
@@
In this study, we propose a clustering-based variable grouping scheme in sub-QUBO formalism. Firstly, we construct a correlation matrix to capture the joint contributions of variables to the objective function. The correlation matrix can be derived from the QUBO matrix and the current solution. Then, we employ the multi-view spectral clustering method [25, 26] to group variables into clusters according to their correlations. We evaluate the performance of our algorithm on Max-Cut problems of randomly generated regular graphs and Erdős–Rényi graphs. Our numerical simulations demonstrate that our method consistently achieves lower objective function values than existing approaches.
@@ II.1 QUBO Problems and QAOA
where $\mathbf{x}$ is a vector of binary variables, $Q$ is a square matrix representing the coefficients of the quadratic terms, and $\mathbf{c}$ is a vector containing the coefficients of the linear terms. Given the binary nature of the variables ($x_{i}\in\{0,1\}$), the linear terms can be equivalently transformed into quadratic terms, enabling the entire problem to be reformulated in a purely quadratic form. The significance of the QUBO framework lies in its ability to encompass a broad range of optimization problems. Many COPs of practical significance can be reformulated by encoding constraints and objectives into binary variables and a quadratic cost matrix [27].
@@
Solving the ground state energy of the Ising model is a NP-hard problem [1, 2]. Classical heuristic algorithms can provide approximate solutions [3, 4, 5, 6]. Specialized computers, such as coherent Ising machines, offer alternative methods that can yield accurate approximations at high speeds [31, 11]. Additionally, quantum computing provides another promising approach for tackling Ising problems, as demonstrated in recent experiments using both analog [32, 33, 34, 35, 36] and digital quantum simulations [37, 38, 39, 40].
QAOA [12, 13] is a variational quantum algorithm that approximates the ground state of a Hamiltonian using parameterized quantum circuits. As illustrated in Fig. 1(f), the QAOA involves two key components: the cost Hamiltonian $H_{C}$, which corresponds to the Ising Hamiltonian in our case, and the mixer Hamiltonian $H_{M}=\sum_{i}\sigma_{x}^{i}$ that facilitates exploration of the solution space.
@@
Despite growing interest, the experimental implementation of QAOA remains constrained by hardware limitations. Recent analog realizations using neutral-atom arrays have scaled to hundreds of qubits and demonstrated promising results on combinatorial problems [41], while gate-based implementations on superconducting and trapped-ion devices are typically limited to fewer than 100 qubits and shallow circuit depths ($p\leq 3$) [16]. These constraints significantly limit the problem sizes that can be realistically addressed. As a result, QAOA has yet to demonstrate consistent performance advantages over classical heuristics in practical settings [42]. These limitations motivate hybrid frameworks that leverage both classical solvers and quantum subroutines to extend QAOA’s applicability.
@@ II.2 The Sub-QUBO Formalism
In this formalism, three components allow flexible implementation choices. The first is the grouping method, discussed in detail in Sec. II.3. The second component is the sub-QUBO solver, which can be quantum annealing, QAOA, or classical algorithms. In this work, we focus primarily on QAOA. Finally, the classical search that forces the solution to a local minimum can also switch between multiple choices. In our work, we employ a simple greedy local search to emphasize the role of the quantum subroutine, while previous studies have explored other classical search methods, including the tabu search [6, 43] as used in Ref. [19]. A discussion on the performance of tabu search within this formalism is provided in appendix A.
@@ II.3 Grouping Methods in the Sub-QUBO Formalism
The effectiveness of hybrid annealing algorithms critically depends on how sub-QUBOs are extracted. Ref. [18] introduced the hybrid annealing framework and several grouping strategies, including random selection and an Impact-Indexing approach (originally referred to as gains-based). The Impact-Indexing method measures the ’impact’ of each variable by evaluating how much the objective function changes when that variable is flipped. Initially, Ref. [18] formed sub-QUBOs by choosing the variables with the largest negative impacts.
@@ IV.1 Dataset
IV.2 Parameterization of the Algorithm
As shown in Fig. 2, the sub-QUBO framework integrates a classical subroutine, a grouping method, and a quantum subroutine to address the challenges of solving QUBO problems. The classical subroutine quickly converges to local minima but may get stuck in suboptimal regions, while the quantum subroutine explores the broader solution space probabilistically. The grouping method bridges the two by partitioning variables into smaller batches for quantum processing, enhancing both optimization quality and efficiency.
For the quantum subroutine, we employ a single-layer Quantum Approximate Optimization Algorithm (QAOA), as it aligns well with the current capabilities of available quantum hardware. The single-layer QAOA circuit consists of three components: state preparation, Hamiltonian evolution of the target Ising model, and Hamiltonian evolution of the mixer Hamiltonian. The state preparation involves applying Hadamard gates to all qubits, which can be implemented efficiently. The QUBO Hamiltonian for MaxCut problems is d-sparse, where $d=\max_{v\in V}(\text{deg}(v))+1$, and the mixer Hamiltonian is 1-local. These structures allow for efficient Hamiltonian simulation [62, 63].
Our implementation, based on Qiskit [64], simulates the quantum subroutine using 1024 shots per measurement to evaluate observables. The parameter optimization is conducted using the gradient-free COBYLA algorithm [65], chosen for its robustness in handling noisy, high-dimensional landscapes. The COBYLA optimizer is initialized with a trust region radius of $1$ at the beginning of the optimization, and gradually reduces this radius to $10^{-4}$ at termination, ensuring precise convergence to the optimal parameter set. Typically, each QAOA execution requires approximately 20 to 30 measurements to converge. Our complete simulation code is openly available under an Apache-2.0 licence at [66].
The classical subroutine can utilize a variety of local search methods, such as greedy search or tabu search. In this study, we employ a single-bit-flip greedy descent: starting from an initial binary vector, we scan sequentially over all variables, compute the energy change
$\Delta E_{i}=(1-2x_{i})\Bigl{(}Q_{ii}+2\sum_{j\neq i}Q_{ij}x_{j}\Bigr{)},$ (12)
flip $x_{i}$ whenever $\Delta E_{i}<0$, and repeat the sweep until no improving move is found. The resulting routine converges deterministically to a local optimum and requires at most $O(n^{2})$ arithmetic operations per sweep. Despite its simplicity, this method has demonstrated satisfactory performance when combined with the quantum subroutine, underscoring the practical potential of this hybrid framework. A more detailed analysis of using tabu search in this context is provided in appendix A.
IV.3 Performance Evaluation
@@
Figure 4: Performance comparison and computational resource requirements for three sub-QUBO algorithms solving Max-Cut problems with sub-problem sizes ranging from 10 to 24 qubits. The algorithms use Impact-Indexing (red), Certainty-Degree (green), and clustering (blue) as grouping methods. (a,d) Minimum objective function values for 100 randomly generated Max-Cut problems based on (a) 100-node 3-regular graphs and (d) 100-node Erdős–Rényi graphs. Lower values indicate better performance. (b,e) Number of quantum QAOA subroutine calls versus sub-problem size for (b) 3-regular graphs and (e) Erdős–Rényi graphs. (c,f) Total number of QAOA iterations versus sub-problem size for (c) 3-regular graphs and (f) Erdős–Rényi graphs. Shaded areas represent 95% confidence intervals.
@@
In Fig. 4(b) and (e), we observe a sharp decrease in the number of quantum subroutine calls when the sub-QUBO size increases from $15$ to $16$ for the Certainty-Degree and Impact-Indexing grouping methods. Further numerical experiments on varying Max-Cut problem sizes using three grouping methods reveal that both Certainty-Degree and Impact-Indexing methods exhibit sudden drops in the number of subroutine calls at specific ratios of sub-QUBO size to problem size. In contrast, the clustering method demonstrates a more consistent behavior, with the number of quantum subroutine calls approximately inversely proportional to the sub-QUBO size across all problem sizes. Detailed discussions of these results are conducted in appendix B.
@@ Appendix B Algorithm Complexity Analysis
For a QUBO or Max-Cut instance on an $n$-vertex graph with $m$ edges, a $p$-layer QAOA circuit applies one entangling gate per edge and one single-qubit rotation per qubit in each layer, leading to a total gate count of $O(pm)$ and a circuit depth of $O(p\Delta)$ on hardware with maximum degree $\Delta$ [12, 67]. On fully connected quantum devices, where entangling gates can be applied in parallel, this depth reduces to $O(p)$ [68].
The cost of evaluating a single parameter vector to additive precision $\varepsilon$ requires $O(1/\varepsilon^{2})$ measurements due to Hoeffding’s bound [69]. Given that the outer-loop optimizer requires $N_{\mathrm{eval}}$ such evaluations, where heuristic methods typically keep $N_{\mathrm{eval}}$ polynomial for fixed $p$ [70, 71], the total runtime depends critically on the number of these evaluations. Additionally, in the large system limit, analytical expressions can be evaluated in $O(16^{p})$ time, reflecting the exponential scaling with $p$ [72].
With a fixed $p$, the quantum subroutine’s time complexity remains effectively constant, and the primary factor influencing overall runtime becomes the number of quantum subroutine calls. The following section examines how this call count scales with the problem size $N$ and the sub-QUBO size $d$ under different grouping strategies, offering insights into the practical scalability of the approach.